# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ 4.2 $2$-soliton solution
$u(x,t)=12\frac{3+4\cosh(2x-8t)+\cosh(4x-64t)}{(3\cosh(x-28t)+\cosh(3x-36t))^{2}},$ (76)
@@
That is, for sufficiently large time, the solution is dominated by a larger soliton moving with speed 16 and a smaller soliton moving with speed 4. To compare with our small-time analysis in section 2, we note that the initial condition (77) has poles at $x_{0}=\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}$ and is a member of (4) with $A_{0}=-6$.
Another way of writing (76) is
@@
$s(t)\sim\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}-2^{2/3}(3t)^{1/3}+12t,\quad\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}+2^{2/3}\left(\frac{1}{2}\pm\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3}+12t$ (78)
@@
The $\mathcal{O}(t^{1/3})$ part of (78) agrees with our prediction (57), that comes from our inner problem for the case $A_{0}=-6$ (and involves the rational solution of PII for $\alpha=2$). The $\mathcal{O}(t)$ terms in (78) are all $12t$; this result agrees with our prediction (118) from appendix B.3 (found by considering the effect of higher-order terms in (3)). Further, these $\mathcal{O}(t)$ terms in (78) are consistent with the well-known results that, for the 2-soliton solution, there are three singularities that initially move out of each of $x_{0}=\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}$ in three equally spaced directions but then all begin to move to the right. More specifically, taking the closest $x_{0}$ to the real-$x$ axis as an example ($n=0$), we can label
@@
Figure 13: Solution profiles and phase portraits for (a) the $2$-soliton solution (76), which evolves from $u(x,0)=6\,\mathrm{sech}^{2}x$ and (b) the $3$-soliton solution (80), which evolves from $u(x,0)=12\,\mathrm{sech}^{2}x$.
Another check on our 2-soliton solution is to put $x={\textstyle\frac{1}{2}}\pi\mathrm{i}+(3t)^{1/3}\xi$ into (76) and then take the limit $t\rightarrow 0^{+}$. The result is that
@@
The leading-order part here agrees with the rational solution computed for $\alpha=2$, namely (53), as expected. The correction term can be checked by showing it is a solution of the appropriate ode problem (115) with (102), the details of which are discussed in appendix B.3. In summary, by carefully studying the exact solution for the $2$-soliton solution we are able to check our matched asymptotics, including the correction terms, providing additional confidence that our analysis is correct.
@@ 4.3 $3$-soliton solution
(see [75]). This solution involves of three solitons moving to the right, with the long-time limit
@@
$s_{3}(t)\sim{\textstyle\frac{1}{2}}\pi\mathrm{i}+(-10+6\sqrt{5})^{1/3}\left(-\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3}+24t,\quad s_{4}(t)\sim{\textstyle\frac{1}{2}}\pi\mathrm{i}+(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3}+24t,$
@@
To compare these results for the $3$-soliton solution with our analysis in section 2, we note that (79) is a member of (4) with $A_{0}=-12$. Thus, the relevant leading-order results are (59)–(61), which come from the roots of $f_{0}$ when $\alpha=3$ (that are derived via the rational solution of PII for $\alpha=3$). The $\mathcal{O}(t^{1/3})$ parts of $s_{1}(t)$ to $s_{6}(t)$ above agree with (59)–(61), as expected, but nevertheless this comparison acts as an important check on our leading-order asymptotic analysis. Further, the $\mathcal{O}(t)$ terms in $s_{1}(t)$ to $s_{6}(t)$ above are all $24t$. These terms agree with the prediction (119) that comes from considering higher-order effects in appendix B.3.
As an additional check, we set $x={\textstyle\frac{1}{2}}\pi\mathrm{i}+(3t)^{1/3}\xi$ in (80) and take the limit $t\rightarrow 0^{+}$, to give
@@
as $t\rightarrow 0^{+}$. The leading-order part of this expression agrees with the rational solution of $f_{0}$ for $\alpha=3$, given by (54), as required. Further, the correction term can be shown to satisfy the appropriate problem (115) with (102), as mentioned in appendix B.3. These tests, including those for higher-order terms, again provide confidence that our matched asymptotics is correct.
@@ 5 Discussion
Solutions of the KdV equation (1) are well known to exhibit dispersive waves that travel in the negative-$x$ direction. For initial conditions that are analytic functions of $x$, we expect the amplitude of these waves to be exponentially small in time. Therefore, they cannot be captured by a naive algebraic expansion in powers of $t$, but instead can be described by considering exponentially-small corrections to the algebraic series and employing Stokes phenomenon in complex-$x$ plane to observe where the waves appear on the real-$x$ axis. One of the goals of our study is to derive the asymptotic form of these waves in the small-time limit and establish connections between this real-valued behaviour and certain singularity structures in the complex-$x$ plane.
@@
For the purposes of our study, a key observation is that (the analytic continuation of) solutions of the KdV equation (1) have singularities in the complex plane that are all double poles with a strength 2-2 (i.e., (2)). Therefore, we have been motivated to study pole dynamics for the KdV equation with a class of initial conditions that also have double poles, each with a strength $A_{0}$, where in general $A_{0}\neq-2$ (see (2)). By applying techniques in exponential asymptotics in the complex plane, we show that dispersive waves behave like (18) on the real line in the small-time limit (see also (5)). The dependence of (18) on $A_{0}$ and $x_{0}$ demonstrates how the strength and location of the double-pole singularities of the analytic continuation of the initial condition have an explicit influence on crucial real-valued wave-like behaviour. Further, without any reference to inverse-scattering techniques or the Painlevé II equation, our exponential asymptotics is able to identify the exceptional nature of the triangular values $A_{0}=-N(N+1)$ in the $\mathrm{sech}^{2}$-type initial conditions (4). More generally, these asymptotic results highlight the role that applied complex analysis can play in analysing real-valued pde solutions.
@@
Next, we have used matched asymptotic expansions to describe how the double poles of KdV emerge from the double poles of the initial condition, and how they propagate for small time. In the neighbourhood of a double pole of the initial condition, $x=x_{0}\notin\mathbb{R}$, we show that the early-time dynamics is driven by a solution of the PII equation (36), written in terms of the similarity variable $\xi=(x-x_{0})/(3t)^{1/3}$, whose constant $\alpha$ is related to $A_{0}$ via (38). This PII problem comes from writing $u=(3t)^{-2/3}f(\xi,t)$, taking the leading-order term $f\sim f_{0}(\xi)$ as $t\rightarrow 0^{+}$, and making the change of variable (31). While there are well-known links between KdV (and modified KdV) and PII, for example via large-time asymptotics or the small-dispersion limit, these typically involve tronquée solutions of the homogeneous version of PII (i.e., with $\alpha=0$) that are real for real values of the independent variable [19, 48, 73], whereas our situation is different because it involves (decreasing) tri tronquée solutions of the inhomogeneous version of PII (i.e., with $\alpha\neq 0$) that is complex on the real line.
@@
Based on the inner problem centred on each $x=x_{0}$, we conjecture that, generically, an infinite number of double poles emerge from $x=x_{0}$. Each one of these poles initially moves like $s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}$ as $t\rightarrow 0^{+}$, where $\xi_{0}$ is one of the poles of PII with residue +1+1 (the poles of PII with residue 1-1 do not correspond to double poles of KdV). Given the tritronquée structure, this implies that there is a lattice of such poles in the $\xi$ plane, bounded by two strings of poles that tend towards the anti-Stokes lines at $\theta=\pi/3$ and $\theta=-\pi$. It appears that the poles that line up close to $\theta=-\pi$ will have the strongest effect on the dispersive waves in the original real-valued problem (1). We can, in principle, compute more terms for the location of each singularity $x=s(t)$ in the small-time limit by considering higher-order correction terms $f\sim f_{0}(\xi)+(3t)^{1/3}f_{1}(\xi)+(3t)^{2/3}f_{2}(\xi)$ to give $s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+(3t)^{2/3}\xi_{1}+(3t)\xi_{2}$ as $t\rightarrow 0^{+}$, where $\xi_{1}$ comes from solving the linear (but complicated) problem for $f_{1}$, $\xi_{2}$ comes from the (even more complicated) problem for $f_{2}$, and so on. The problems for $f_{1}$, $f_{2}$, $\ldots$, depend on the parameters $A_{1}$, $A_{2}$, $\ldots$, in (21), which ultimately links all of these small-time results to the local expansion of the singularity of the initial condition (21).
@@
We have applied some formal techniques in transseries asymptotics to approximate the location of the string of singularities for our PII problem that tend to the negative real $\xi$ axis as $|\xi|\rightarrow\infty$. To leading order, these results imply that corresponding poles of KdV move like $s(t)\sim x_{0}-3\pi^{2/3}n^{2/3}t^{1/3}$ as $t\rightarrow 0^{+}$, $n\rightarrow\infty$, where $n$ is a natural number. On the other hand, our exponential asymptotics suggest that the crests of the dispersive waves also move like $\mathrm{Re}(x_{0})-3\pi^{2/3}n^{2/3}t^{1/3}$ on the real line (see (19)). Clearly, there is a link between certain complex singularities of a solution of the KdV equation and the wavelike behaviour of that solution on the real line.
@@
We have identified a number of special cases. For example, when the parameter $\alpha=N$ for $N\in\mathbb{N}$, our tritronquée solutions of PII, with an infinite number of poles, reduce to rational solutions, which have a finite number of poles. These solutions are relevant for $\mathrm{sech}^{2}$-type initial conditions (4), where the prefactor $A_{0}$ is a triangular number $N(N+1)$. We have shown that for $2$- and $3$-soliton solutions, our asymptotics agree with the exact solutions. Other special cases come from $A_{0}=-2$, for which the singularities move with a different temporal scaling. For example, if an initial condition has a double pole at $x=x_{0}$ with local behaviour (21), where $A_{0}=-2$ but $A_{1}\neq 0$, then we expect a singularity to emerge from $x_{0}$ with the scaling $s(t)=x_{0}+\mathcal{O}(t^{2/3})$ (see (126)). Further, if $A_{0}=-2$, $A_{1}=0$, $A_{2}\neq 0$, then the scaling changes to $s(t)=x_{0}+\mathcal{O}(t)$ (see (127)).
We regard our work as a preliminary study that has left a number of unresolved issues and suggests numerous open problems worth attention. We list some of these here in the remainder of this section.
@@
  • The exponential asymptotics detailed in subsection 2.2 and appendix A provides leading order behaviour of the dispersive waves for the generic case $A_{0}\neq-N(N+1)$ as well as the special case $A_{0}=-N(N+1)$, the latter depending on the coefficient of the correction term in (21), namely $A_{1}$. It remains to derive the correction terms for the generic case $A_{0}\neq-N(N+1)$, which will involve both $A_{0}$ and $A_{1}$. This type of higher-order analysis for this nonlinear problem within the context of exponential asymptotics is likely to be nontrivial.
@@
    which is not surprising as dispersive waves are exponentially small in the small-time limit and therefore the nonlinear term in the full KdV equation does not play a role in deriving the form of the late-order terms (15) nor the exponentially-small contributions (16). The matching into the inner region near $x=x_{0}$ will be different for the linear problem, however, since the function $f_{0}$ will satisfy the linear ode $2f_{0}-\xi f_{0}^{\prime}+f^{\prime\prime\prime}=0$ instead of (27). Therefore, applying the series (88), the recurrence relation (90) will be $a_{n}=(3n-1)(3n+1)a_{n-1}$, which can be solved easily to evaluate the limit (89), suggesting $\Lambda=\mathrm{i}A_{0}/3^{3/4}\sqrt{\pi}$ and
    $U_{\mbox{\footnotesize dis}}\sim-\frac{2\sqrt{\pi}A_{0}}{3^{3/4}}\frac{(-x)^{1/4}\,\mathrm{e}^{-y_{0}(-x/3t)^{1/2}}}{t^{3/4}}\,\cos\left(\frac{2(-x)^{3/2}}{3(3t)^{1/2}}-\frac{\pi}{4}\right)$ (81)
@@
    as $x\rightarrow$-$\infty$, $t\rightarrow 0^{+}$ (cf. (5)). The claim (81) could be checked using integral transforms and appropriate limits and then expanded upon to hold for different initial conditions other than those with the property (3).
@@
  • For the well-studied initial conditions of the form (4), it remains to explain how the singularity structure for (43)–(45) changes as the parameter $A_{0}$ varies from slightly greater than $A_{0}=-N(N+1)$ for a natural number NN ($\alpha<N$) to slightly less than N(N+1)-N(N+1) ($\alpha>N$), and how these changes affect qualitative behaviour for KdV on the real line. For example, in figure 14 we show three phase portraits illustrating solutions of (43)–(45) for $\alpha=2.99$, $3$ and $3.01$. It is not clear how the lattice of infinitely many poles are arranged in the two limits $\alpha\rightarrow 3^{-}$ and $\alpha\rightarrow 3^{+}$, given only nine are present for the borderline case $\alpha=3$.
@@
    Furthermore, with the special initial condition (4), when $A_{0}$ is slightly greater than 12-12 ($\alpha\lesssim 3$), there will be two solitons that eventually evolve to the right, while for $A_{0}=-12$ ($\alpha=3$) or slightly less than 12-12 ($\alpha\gtrsim 3$), there will be three. Therefore, while we can speculate that the pole dynamics for the solitons when $A_{0}$ is slightly less than 12-12 is qualitatively similar to that for $A_{0}=-12$ (discussed in subsection 4.3), it is not at all clear what happens for $A_{0}$ slightly greater than 12-12.
@@
    Figure 14: Phase portraits of numerical solutions of (43)–(45) computed for $\alpha=2.99$ (left), $3$ (centre) and $3.01$ (right).
@@
  • For initial conditions with singularities of the form (3), we have not explored the parameter range $A_{0}>1/4$, which corresponds to complex values of $\alpha$. This parameter range would be relevant for “dip”-type versions of (4) or (6). It would be interesting to study relevant solutions of PII, which may behave differently to the cases in which $\alpha$ is real-valued.
@@
  • While the asymptotic methods used in subsection 3.6 are able to estimate the location of the important singularities of (43)–(45) that align near the negative real $\xi$ axis as $|\xi|\rightarrow\infty$ (see (73)–(74)), a more refined version of these techniques, using a two-parameter transseries [5, 72], could be developed to describe the full singularity structure between the two active anti-Stokes lines (the negative real $\xi$ axis and $\mathrm{arg}(\xi)=\pi/3$).
@@
    $u_{0}\sim\frac{A}{(x-x_{0})^{\beta}}\quad\mbox{as}\quad x\rightarrow x_{0},$ (82)
    with $\beta\neq 2$, the details for the matched asymptotic expansions must be very different to the case we have considered (namely $\beta=2$). For a start, the second equation in (9) must be replaced by
    $u_{1}\sim\frac{6\beta A^{2}}{(x-x_{0})^{2\beta+1}}+\frac{\beta(\beta+1)(\beta+2)A}{(x-x_{0})^{\beta+3}}\quad\mbox{as}\quad x\rightarrow x_{0}.$ (83)
    For $\beta>2$, the first term on the right-hand side of (83) dominates the second, and thus the inner problem must have
@@
    (cf. (23) and (25)). It turns out that the third-order dispersive term does not appear in the leading-order problem, governed by
    $\frac{3\beta}{\beta+1}f_{0}-\frac{3}{\beta+1}\xi f_{0}^{\prime}+6f_{0}f_{0}^{\prime}=0$
    (cf. (27)), which suggests a further rescaling will be required in order to bring in the as-yet neglected higher-order term. For $0<\beta<2$, the second term on the right-hand side of (83) dominates the first, and so the inner problem has
@@
    $\beta f_{0}-\xi f_{0}^{\prime}+f_{0}^{\prime\prime\prime}=0,$

### v2

@@ 4.2 $2$-soliton solution
$u(x,t)=12\frac{3+4\cosh(2x-8t)+\cosh(4x-64t)}{(3\cosh(x-28t)+\cosh(3x-36t))^{2}},$ (4.2)
@@
That is, for sufficiently large time, the solution is dominated by a larger soliton moving with speed 16 and a smaller soliton moving with speed 4. To compare with our small-time analysis in section 2, we note that the initial condition (4.3) has poles at $x_{0}=\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}$ and is a member of (1.4) with $A_{0}=-6$.
Another way of writing (4.2) is
@@
$s(t)\sim\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}-2^{2/3}(3t)^{1/3}+12t,\quad\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}+2^{2/3}\left(\frac{1}{2}\pm\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3}+12t$ (4.4)
@@
The $\mathcal{O}(t^{1/3})$ part of (4.4) agrees with our prediction (3.38), that comes from our inner problem for the case $A_{0}=-6$ (and involves the rational solution of PII for $\alpha=2$). The $\mathcal{O}(t)$ terms in (4.4) are all $12t$; this result agrees with our prediction (B.25) from appendix B.3 (found by considering the effect of higher-order terms in (1.3)). Further, these $\mathcal{O}(t)$ terms in (4.4) are consistent with the well-known results that, for the 2-soliton solution, there are three singularities that initially move out of each of $x_{0}=\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}$ in three equally spaced directions but then all begin to move to the right. More specifically, taking the closest $x_{0}$ to the real-$x$ axis as an example ($n=0$), we can label
@@
Figure 13: Solution profiles and phase portraits for (a) the $2$-soliton solution (4.2), which evolves from $u(x,0)=6\,\mathrm{sech}^{2}x$ and (b) the $3$-soliton solution (4.6), which evolves from $u(x,0)=12\,\mathrm{sech}^{2}x$.
Another check on our 2-soliton solution is to put $x={\textstyle\frac{1}{2}}\pi\mathrm{i}+(3t)^{1/3}\xi$ into (4.2) and then take the limit $t\rightarrow 0^{+}$. The result is that
@@
The leading-order part here agrees with the rational solution computed for $\alpha=2$, namely (3.34), as expected. The correction term can be checked by showing it is a solution of the appropriate ode problem (B.22) with (B.9), the details of which are discussed in appendix B.3. In summary, by carefully studying the exact solution for the $2$-soliton solution we are able to check our matched asymptotics, including the correction terms, providing additional confidence that our analysis is correct.
@@ 4.3 $3$-soliton solution
(see [85]). This solution involves three solitons moving to the right, with the long-time limit
@@
$s_{3}(t)\sim{\textstyle\frac{1}{2}}\pi\mathrm{i}+(-10+6\sqrt{5})^{1/3}\left(-\frac{1}{2}+\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3}+24t,\,\,s_{4}(t)\sim{\textstyle\frac{1}{2}}\pi\mathrm{i}+(10+6\sqrt{5})^{1/3}\left(\frac{1}{2}-\frac{\sqrt{3}}{2}\mathrm{i}\right)(3t)^{1/3}+24t,$
@@
To compare these results for the $3$-soliton solution with our analysis in section 2, we note that (4.5) is a member of (1.4) with $A_{0}=-12$. Thus, the relevant leading-order results are (3.40)–(3.42), which come from the roots of $f_{0}$ when $\alpha=3$ (that are derived via the rational solution of PII for $\alpha=3$). The $\mathcal{O}(t^{1/3})$ parts of $s_{1}(t)$ to $s_{6}(t)$ above agree with (3.40)–(3.42), as expected, but nevertheless this comparison acts as an important check on our leading-order asymptotic analysis. Further, the $\mathcal{O}(t)$ terms in $s_{1}(t)$ to $s_{6}(t)$ above are all $24t$. These terms agree with the prediction (B.26) that comes from considering higher-order effects in appendix B.3.
As an additional check, we set $x={\textstyle\frac{1}{2}}\pi\mathrm{i}+(3t)^{1/3}\xi$ in (4.6) and take the limit $t\rightarrow 0^{+}$, to give
@@
as $t\rightarrow 0^{+}$. The leading-order part of this expression agrees with the rational solution of $f_{0}$ for $\alpha=3$, given by (3.35), as required. Further, the correction term can be shown to satisfy the appropriate problem (B.22) with (B.9), as mentioned in appendix B.3. These tests, including those for higher-order terms, again provide confidence that our matched asymptotics is correct.
@@ 5 Discussion
Solutions of the KdV equation (1.1) are well known to exhibit dispersive waves that travel in the negative-$x$ direction. For initial conditions that are analytic functions of $x$, we expect the amplitude of these waves to be exponentially small in time. Therefore, they cannot be captured by a naive algebraic expansion in powers of $t$, but instead can be described by considering exponentially-small corrections to the algebraic series and employing Stokes phenomenon in complex-$x$ plane to observe where the waves appear on the real-$x$ axis. One of the goals of our study is to derive the asymptotic form of these waves in the small-time limit and establish connections between this real-valued behaviour and certain singularity structures in the complex-$x$ plane.
@@
For the purposes of our study, a key observation is that (the analytic continuation of) solutions of the KdV equation (1.1) have singularities in the complex plane that are all double poles with a strength 2-2 (i.e., (1.2)). Therefore, we have been motivated to study pole dynamics for the KdV equation with a class of initial conditions that also have double poles, each with a strength $A_{0}$, where in general $A_{0}\neq-2$ (see (1.2)). By applying techniques in exponential asymptotics in the complex plane, we show that dispersive waves behave like (2.12) on the real line in the small-time limit (see also (1.5)). The dependence of (2.12) on $A_{0}$ and $x_{0}$ demonstrates how the strength and location of the double-pole singularities of the analytic continuation of the initial condition have an explicit influence on crucial real-valued wave-like behaviour. Further, without any reference to inverse-scattering techniques or the Painlevé II equation, our exponential asymptotics is able to identify the exceptional nature of the triangular values $A_{0}=-N(N+1)$ in the $\mathrm{sech}^{2}$-type initial conditions (1.4). More generally, these asymptotic results highlight the role that applied complex analysis can play in analysing real-valued pde solutions.
@@
Next, we have used matched asymptotic expansions to describe how the double poles of KdV emerge from the double poles of the initial condition, and how they propagate for small time. In the neighbourhood of a double pole of the initial condition, $x=x_{0}\notin\mathbb{R}$, we show that the early-time dynamics is driven by a solution of the PII equation (3.14), written in terms of the similarity variable $\xi=(x-x_{0})/(3t)^{1/3}$, whose constant $\alpha$ is related to $A_{0}$ via (3.16). This PII problem comes from writing $u=(3t)^{-2/3}f(\xi,t)$, taking the leading-order term $f\sim f_{0}(\xi)$ as $t\rightarrow 0^{+}$, and making the change of variable (3.9). While there are well-known links between KdV (and modified KdV) and PII, for example via large-time asymptotics or the small-dispersion limit, these typically involve tronquée solutions of the homogeneous version of PII (i.e., with $\alpha=0$) that are real for real values of the independent variable [22, 54, 83], whereas our situation is different because it involves (decreasing) tri tronquée solutions of the inhomogeneous version of PII (i.e., with $\alpha\neq 0$) that is complex on the real line.
@@
Based on the inner problem centred on each $x=x_{0}$, we conjecture that, generically, an infinite number of double poles emerge from $x=x_{0}$. Each one of these poles initially moves like $s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}$ as $t\rightarrow 0^{+}$, where $\xi_{0}$ is one of the poles of PII with residue +1+1 (the poles of PII with residue 1-1 do not correspond to double poles of KdV). Given the tritronquée structure, this implies that there is a lattice of such poles in the $\xi$ plane, bounded by two strings of poles that tend towards the anti-Stokes lines at $\theta=\pi/3$ and $\theta=-\pi$. It appears that the poles that line up close to $\theta=-\pi$ will have the strongest effect on the dispersive waves in the original real-valued problem (1.1). We can, in principle, compute more terms for the location of each singularity $x=s(t)$ in the small-time limit by considering higher-order correction terms $f\sim f_{0}(\xi)+(3t)^{1/3}f_{1}(\xi)+(3t)^{2/3}f_{2}(\xi)$ to give $s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+(3t)^{2/3}\xi_{1}+(3t)\xi_{2}$ as $t\rightarrow 0^{+}$, where $\xi_{1}$ comes from solving the linear (but complicated) problem for $f_{1}$, $\xi_{2}$ comes from the (even more complicated) problem for $f_{2}$, and so on. The problems for $f_{1}$, $f_{2}$, $\ldots$, depend on the parameters $A_{1}$, $A_{2}$, $\ldots$, in (2.15), which ultimately links all of these small-time results to the local expansion of the singularity of the initial condition (2.15).
@@
We have applied some formal techniques in transseries asymptotics to approximate the location of the string of singularities for our PII problem that tend to the negative real $\xi$ axis as $|\xi|\rightarrow\infty$. To leading order, these results imply that corresponding poles of KdV move like $s(t)\sim x_{0}-3\pi^{2/3}n^{2/3}t^{1/3}$ as $t\rightarrow 0^{+}$, $n\rightarrow\infty$, where $n$ is a natural number. On the other hand, our exponential asymptotics suggest that the crests of the dispersive waves also move like $\mathrm{Re}(x_{0})-3\pi^{2/3}n^{2/3}t^{1/3}$ on the real line (see (2.13)). Clearly, there is a link between certain complex singularities of a solution of the KdV equation and the wavelike behaviour of that solution on the real line.
@@
We have identified a number of special cases. For example, when the parameter $\alpha=N$ for $N\in\mathbb{N}$, our tritronquée solutions of PII, with an infinite number of poles, reduce to rational solutions, which have a finite number of poles. These solutions are relevant for $\mathrm{sech}^{2}$-type initial conditions (1.4), where the prefactor $A_{0}$ is a triangular number $N(N+1)$. We have shown that for $2$- and $3$-soliton solutions, our asymptotics agree with the exact solutions. Other special cases come from $A_{0}=-2$, for which the singularities move with a different temporal scaling. For example, if an initial condition has a double pole at $x=x_{0}$ with local behaviour (2.15), where $A_{0}=-2$ but $A_{1}\neq 0$, then we expect a singularity to emerge from $x_{0}$ with the scaling $s(t)=x_{0}+\mathcal{O}(t^{2/3})$ (see (B.33)). Further, if $A_{0}=-2$, $A_{1}=0$, $A_{2}\neq 0$, then the scaling changes to $s(t)=x_{0}+\mathcal{O}(t)$ (see (B.34)).
@@
We regard our work as a preliminary study that has left a number of unresolved issues and suggests numerous open problems worth attention. We list some of these here in the remainder of this section. The exponential asymptotics detailed in subsection 2.2 and appendix A provides leading order behaviour of the dispersive waves for the generic case $A_{0}\neq-N(N+1)$ as well as the special case $A_{0}=-N(N+1)$, the latter depending on the coefficient of the correction term in (2.15), namely $A_{1}$. It remains to derive the correction terms for the generic case $A_{0}\neq-N(N+1)$, which will involve both $A_{0}$ and $A_{1}$. This type of higher-order analysis for this nonlinear problem within the context of exponential asymptotics is likely to be nontrivial.
@@
which is not surprising as dispersive waves are exponentially small in the small-time limit and therefore the nonlinear term in the full KdV equation does not play a role in deriving the form of the late-order terms (2.9) nor the exponentially-small contributions (2.10). The matching into the inner region near $x=x_{0}$ will be different for the linear problem, however, since the function $f_{0}$ will satisfy the linear ode
$2f_{0}-\xi\frac{\mathrm{d}f_{0}}{\mathrm{d}\xi}+\frac{\mathrm{d}^{3}f_{0}}{\mathrm{d}\xi^{3}}=0$
instead of (3.5). Therefore, applying the series (A.5), the recurrence relation (A.7) will be $a_{n}=(3n-1)(3n+1)a_{n-1}$, which can be solved easily to evaluate the limit (A.6), suggesting $\Lambda=\mathrm{i}A_{0}/3^{3/4}\sqrt{\pi}$ and
$U_{\mbox{\footnotesize dis}}\sim-\frac{2\sqrt{\pi}A_{0}}{3^{3/4}}\frac{(-x)^{1/4}\,\mathrm{e}^{-y_{0}(-x/3t)^{1/2}}}{t^{3/4}}\,\cos\left(\frac{2(-x)^{3/2}}{3(3t)^{1/2}}-\frac{\pi}{4}\right)$ (5.1)
@@
as $x\rightarrow$-$\infty$, $t\rightarrow 0^{+}$ (cf. (1.5)). The claim (5.1) could be checked using integral transforms and appropriate limits and then expanded upon to hold for different initial conditions other than those with the property (1.3).
@@
For the well-studied initial conditions of the form (1.4), it remains to explain how the singularity structure for (3.23)–(3.25) changes as the parameter $A_{0}$ varies from slightly greater than $A_{0}=-N(N+1)$ for a natural number NN ($\alpha<N$) to slightly less than N(N+1)-N(N+1) ($\alpha>N$), and how these changes affect qualitative behaviour for KdV on the real line. For example, in figure 14 we show three phase portraits illustrating solutions of (3.23)–(3.25) for $\alpha=2.99$, $3$ and $3.01$. It is not clear how the lattice of infinitely many poles are arranged in the two limits $\alpha\rightarrow 3^{-}$ and $\alpha\rightarrow 3^{+}$, given only nine are present for the borderline case $\alpha=3$.
@@
Furthermore, with the special initial condition (1.4), when $A_{0}$ is slightly greater than 12-12 ($\alpha\lesssim 3$), there will be two solitons that eventually evolve to the right, while for $A_{0}=-12$ ($\alpha=3$) or slightly less than 12-12 ($\alpha\gtrsim 3$), there will be three. Therefore, while we can speculate that the pole dynamics for the solitons when $A_{0}$ is slightly less than 12-12 is qualitatively similar to that for $A_{0}=-12$ (discussed in subsection 4.3), it is not at all clear what happens for $A_{0}$ slightly greater than 12-12.
@@
Figure 14: Phase portraits of numerical solutions of (3.23)–(3.25) computed for $\alpha=2.99$ (left), 33 (centre) and $3.01$ (right).
@@
For initial conditions with singularities of the form (1.3), we have not explored the parameter range $A_{0}>1/4$, which corresponds to complex values of $\alpha$. This parameter range would be relevant for “dip”-type versions of (1.4) or (1.6). It would be interesting to study relevant solutions of PII, which may behave differently to the cases in which $\alpha$ is real-valued.
@@
While the asymptotic methods used in subsection 3.7 are able to estimate the location of the important singularities of (3.23)–(3.25) that align near the negative real $\xi$ axis as $|\xi|\rightarrow\infty$ (see (3.54)–(3.55)), a more refined version of these techniques, using a two-parameter transseries [6, 82], could be developed to describe the full singularity structure between the two active anti-Stokes lines (the negative real $\xi$ axis and $\mathrm{arg}(\xi)=\pi/3$).
@@
$u_{0}\sim\frac{A}{(x-x_{0})^{\beta}}\quad\mbox{as}\quad x\rightarrow x_{0},$ (5.2)
with $\beta\neq 2$, the details for the matched asymptotic expansions must be very different to the case we have considered (namely $\beta=2$). For a start, the second equation in (2.3) must be replaced by
$u_{1}\sim\frac{6\beta A^{2}}{(x-x_{0})^{2\beta+1}}+\frac{\beta(\beta+1)(\beta+2)A}{(x-x_{0})^{\beta+3}}\quad\mbox{as}\quad x\rightarrow x_{0}.$ (5.3)
For $\beta>2$, the first term on the right-hand side of (5.3) dominates the second, and thus the inner problem must have
@@
(cf. (3.1) and (3.3)). It turns out that the third-order dispersive term does not appear in the leading-order problem, governed by
$\frac{3\beta}{\beta+1}f_{0}-\frac{3}{\beta+1}\xi\frac{\mathrm{d}f_{0}}{\mathrm{d}\xi}+6f_{0}\frac{\mathrm{d}f_{0}}{\mathrm{d}\xi}=0$
(cf. (3.5)), which suggests a further rescaling will be required in order to bring in the as-yet neglected higher-order term. For $0<\beta<2$, the second term on the right-hand side of (5.3) dominates the first, and so the inner problem has
@@
$\beta f_{0}-\xi\frac{\mathrm{d}f_{0}}{\mathrm{d}\xi}+\frac{\mathrm{d}^{3}f_{0}}{\mathrm{d}\xi^{3}}=0,$